# Filtering for Creativity: Adaptive Prompting for Multilingual Riddle Generation in LLMs

### v1

@@ B.1 Average Token Length Across Pretrained Models
B.2 Average Token Lengths Across Languages
@@
B.3 Cross-Lingual Evaluation of Syntactic Validity
@@
B.4 Average self-BLEU and Distinct-n Pretrained Metrics
@@
Appendix C Appendix: English Riddle Examples
C.1 English Pretrained Riddle Generations
@@
C.2 English Comparison of Fine-Tuned Riddle Generations to Pretrained Counterparts
@@
C.3 English Fine-Tuned Riddles and Their Real-World Counterparts
@@
C.4 English Fine-Tuned Riddle Examples
@@
Appendix D Appendix: Japanese Riddle Examples
D.1 Japanese Pretrained Bilingual Riddle Examples
@@
D.2 Japanese Pretrained vs Fine-Tuned Bilingual Riddle Examples
@@
Zero-Shot 頭はあるが泣くことはない床はあるが寝ることはない口はあるが話すことはないそして、変わるが変わらないものなんだ何なのだろう ”川” (I have a head, but never weep… A River) 羽がなくても空を飛び、目がなくても涙を流すものは何? (“What flies without wings and cries without eyes?”)
@@
D.3 Japanese Fine-Tuned vs Real-World Riddles
@@
Appendix E Appendix: Arabic Riddle Examples
E.1 Arabic Pretrained Bilingual Riddle Examples
@@
E.2 Arabic Pretrained vs Fine-Tuned Bilingual Riddle Examples
@@
E.3 Real-World Riddles vs. Fine-Tuned Arabic Riddles
@@
Appendix F Appendix: Chinese Riddle Examples
F.1 Chinese Pretrained Riddle Examples
@@
F.2 Chinese Fine-Tuned vs Pretrained Riddle Examples
@@
F.3 Chinese Fine-Tuned vs Real-World Riddles
@@ 1 ZH: 口袋里有个圆,白天不见晚上现。
F.4 Chinese Fine-Tuned AOF Examples
@@
Appendix G Appendix: French Riddle Examples
G.1 French Pretrained Riddle Examples
@@
G.2 French Pretrained vs Fine-Tuned
@@
G.3 French Fine-Tuned Riddles and Their Real-World Counterparts
@@ 1 FR: Je vole sans ailes, je pleure sans yeux…
G.4 French Fine-Tuned AOF Examples
@@
Appendix H Appendix:Prompting Methods
H.1 Chinese prompts
@@ H.4 French prompts
Appendix I Appendix: Fined-tuned Training and Evaluation Details
I.1 Dataset Selection and Preparation
We used the BiRdQA dataset Zhang & Wan (2022), a multilingual benchmark designed to test figurative language understanding and commonsense inference. It includes 6,614 English riddles and 8,751 Chinese riddles, each paired with four answer options. Riddles were shuffled at each epoch to prevent memorization, and no synthetic augmentation was applied.
@@ I.2 Training Strategy
I.3 Appendix:Training Set Expansion
@@
I.4 Model Comparison Methodology
I.4.1 Baseline Models
@@ I.4.3 Summary of Findings
I.5 Impact of Multiple-Choice Framing
@@
Appendix J Appendix: AOF Prompt Template and Constraints
@@
  • Novelty Filter: Semantic similarity to known riddles must fall below a threshold ($\theta=0.75$), as measured against BiRdQA Zhang & Wan (2022).
J.1 Semantic Similarity Filtering Equation
@@
J.2 Rejection Sampling Algorithm
@@
J.3 Threshold Sensitivity: Self-BLEU and Distinct-2
@@
Appendix K Appendix: Experimental Configuration Details
@@
  • Adversarial: Distractor-rich prompts based on known LLM vulnerabilities Wallace et al. (2019); Ribeiro et al. (2018).
  • Adaptive Originality Filtering (AOF): Filtering-based prompting for semantic novelty. See Appendix J.
Prompt formatting logic appears in Appendix H
@@
We used BiRdQA Zhang & Wan (2022), which contains:

### v3

@@ E.1 Average Token Length Across Pretrained Models
E.2 Average Token Lengths Across Languages
@@
E.3 Cross-Lingual Evaluation of Syntactic Validity
@@
E.4 Average self-BLEU and Distinct-n Pretrained Metrics
@@
E.5 Average self-BLEU and Distinct-n fined tuned Metrics
Lang. Pair Prompting Method Self-BLEU / Distinct-2
Eng–Arabic Few-Shot 0.233 / 0.826
AOF (Ours) 0.260 / 0.893
Zero-Shot 0.391 / 0.752
Chain-of-Thought 0.326 / 0.831
Adversarial 0.320 / 0.810
Eng–Chinese AOF (Ours) 0.163 / 0.934
Zero-Shot 0.315 / 0.831
Few-Shot 0.349 / 0.787
Chain-of-Thought 0.305 / 0.828
Adversarial 0.400 / 0.757
Eng–Japanese AOF (Ours) 0.177 / 0.915
Zero-Shot 0.431 / 0.752
Few-Shot 0.326 / 0.778
Chain-of-Thought 0.386 / 0.796
Adversarial 0.327 / 0.748
Eng–French AOF (Ours) 0.273 / 0.856
Zero-Shot 0.289 / 0.867
Few-Shot 0.323 / 0.835
Chain-of-Thought 0.256 / 0.892
Adversarial 0.359 / 0.793
Table 9: Self-BLEU (lower is better) and Distinct-2 (higher is better) for fine-tuned GPT-4o across prompting methods. Best combined performance per language pair in bold.
Appendix F Appendix: English Riddle Examples
F.1 English Pretrained Riddle Generations
@@
F.2 English Comparison of Fine-Tuned Riddle Generations to Pretrained Counterparts
@@
F.3 English Fine-Tuned Riddles and Their Real-World Counterparts
@@
F.4 English Fine-Tuned Riddle Examples
@@
Appendix G Appendix: Japanese Riddle Examples
G.1 Japanese Pretrained Bilingual Riddle Examples
@@
G.2 Japanese Pretrained vs Fine-Tuned Bilingual Riddle Examples
@@
Zero-Shot 頭はあるが泣くことはない床はあるが寝ることはない口はあるが話すことはないそして、変わるが変わらないものなんだ何なのだろう "川" (I have a head, but never weep… A River) 羽がなくても空を飛び、目がなくても涙を流すものは何? (“What flies without wings and cries without eyes?”)
@@
G.3 Japanese Fine-Tuned vs Real-World Riddles
@@
Appendix H Appendix: Arabic Riddle Examples
H.1 Arabic Pretrained Bilingual Riddle Examples
@@
H.2 Arabic Pretrained vs Fine-Tuned Bilingual Riddle Examples
@@
H.3 Real-World Riddles vs. Fine-Tuned Arabic Riddles
@@
Appendix I Appendix: Chinese Riddle Examples
I.1 Chinese Pretrained Riddle Examples
@@
I.2 Chinese Fine-Tuned vs Pretrained Riddle Examples
@@
I.3 Chinese Fine-Tuned vs Real-World Riddles
@@ 1 ZH: 口袋里有个圆,白天不见晚上现。
I.4 Chinese Fine-Tuned AOF Examples
@@
Appendix J Appendix: French Riddle Examples
J.1 French Pretrained Riddle Examples
@@
J.2 French Pretrained vs Fine-Tuned
@@
J.3 French Fine-Tuned Riddles and Their Real-World Counterparts
@@ 1 FR: Je vole sans ailes, je pleure sans yeux…
J.4 French Fine-Tuned AOF Examples
@@
Appendix K Appendix:Prompting Methods
K.1 Chinese prompts
@@
Appendix L Appendix: Fined-tuned Training and Evaluation Details
L.1 Dataset Selection and Preparation
We used the BiRdQA dataset (Zhang and Wan, 2022), a multilingual benchmark designed to test figurative language understanding and commonsense inference. It includes 6,614 English riddles and 8,751 Chinese riddles, each paired with four answer options. Riddles were shuffled at each epoch to prevent memorization, and no synthetic augmentation was applied.
@@
L.3 Appendix:Training Set Expansion
@@
L.4 Model Comparison Methodology
L.4.1 Baseline Models
@@
L.5 Impact of Multiple-Choice Framing
@@
Appendix M Appendix: AOF Prompt Template and Constraints
@@
  • Novelty Filter: Semantic similarity to known riddles must fall below a threshold ($\theta=0.75$), as measured against BiRdQA (Zhang and Wan, 2022).
M.1 Semantic Similarity Filtering Equation
@@
M.2 Rejection Sampling Algorithm
@@
M.3 Threshold Sensitivity: Self-BLEU and Distinct-2
@@
Appendix N Appendix: Experimental Configuration Details
@@
  • Adversarial: Distractor-rich prompts based on known LLM vulnerabilities (Wallace et al., 2019; Ribeiro et al., 2018).
  • Adaptive Originality Filtering (AOF): Filtering-based prompting for semantic novelty. See Appendix M.
Prompt formatting logic appears in Appendix K
@@
We used BiRdQA (Zhang and Wan, 2022), which contains: